Let A = [ 0 1 2

1 0 3

4 -3 8 ]

If A is an invertible matrix, then show that.

det (A-1) = 1/det ( A )




1
Expert's answer
2021-05-19T12:57:12-0400

"A^{-1}=\frac{A^T_*}{|A|}"

"|A|=4-2\cdot3=-2"

Minor matrix: each element "m_{ij}" is the determinant of 2x2 matrix, if we remove i-th row and j-th column in matrix "A".

"M=\begin{pmatrix}
 9 & -4&-3 \\
 14 & -8&-4\\
 3&-2&-1 \\
\end{pmatrix}"

Matrix of cofactors: change signs of "m_{12},m_{21},m_{23},m_{32}."

"A_*=\begin{pmatrix}
 9 & 4&-3 \\
 -14 & -8&4\\
 3&2&-1 \\
\end{pmatrix}"

Transpose matrix of cofactors:

"A^T_*=\begin{pmatrix}
 9 & -14&3 \\
 -4 & -8&2\\
 -3&4&1 \\
\end{pmatrix}"

"A^{-1}=-\frac{1}{2}\begin{pmatrix}
 9 & -14&3 \\
 -4 & -8&2\\
 -3&4&1 \\
\end{pmatrix}"

"|A^{-1}|=4.5\cdot4-7\cdot0.5-1.5\cdot10=-0.5=1/|A|"